OpenStudy (anonymous):

How many solutions exist for the system y = 2x + 5 and y = -0.2x - 4? no solution one solution infinitely many solutions

OpenStudy (anonymous):

Set them equal to eachother. 2x+5 = -.2x-4 2.2x = -9 x = -9/2.2 = one real solution.

OpenStudy (anonymous):

ty